OpenStudy (anonymous):

F(x)=xln(x)-x Where x>0, Find F''(x)

OpenStudy (anonymous):

do u need explanation or just answer?

OpenStudy (anonymous):

first for derivative of xlnx you use the product rule let u=x v=lnx u'=1 v'=1/x and derivative of x=1 so \[f'(x)=u'v+v'u\] \[f'(x)=lnx+1 -1\] \[f''(x)=\frac{1}{x}\]

OpenStudy (anonymous):

it should be\[f'(x) = u'v+v'u -1\]